# Clarifying the effect of mean subtraction on Dynamic Mode Decomposition

### v5

@@ Abstract
Pre-processing time-sequential training data by removing the temporal mean has been a point of contention in the Companion matrix formulation of DMD. This stems from the potential of said pre-processing to render DMD equivalent to a temporal Discrete Fourier Transform (DFT). We prove that this equivalence is impossible when the training data is linearly consistent and the order of the DMD approximation exceeds the number of Koopman modes. Since model order and training set size are synonymous in this variant of DMD, the parity of DMD and DFT can, therefore, be indicative of inadequate training data.
@@ 3.1 The Koopman operator
$=~{}\alpha_{1}~{}\underbrace{\psi_{1}\circ\Gamma}~{}+~{}\alpha_{2}~{}\underbrace{\psi_{2}\circ\Gamma}$
@@ 3.2 Dynamic Mode Decomposition
Dynamic Mode Decomposition [20] is an algorithm that takes time-sequential observations and returns estimates of the constituent Koopman eigen-values and modes. The input to DMD is generated by first choosing $m$ observables of interest (collectively referred to as the dictionary), $\{\psi_{i}\}_{i=1}^{m}$, evaluating them along an orbit, $\{\Gamma^{j}\left(\boldsymbol{\mathbf{\upsilon}}_{1}\right)\}_{j=0}^{n}$, and collecting the resulting observations (aka snapshots) as the columns of a matrix $\mathbf{Z}$:
(4) $\boldsymbol{\mathbf{\psi}}:=[\psi_{i}]_{i=1}^{m},\quad\mathbf{Z}~{}:=~{}\left[\boldsymbol{\mathbf{\psi}}\left(\Gamma^{j}\left(\boldsymbol{\mathbf{\upsilon}}_{1}\right)\right)\right]_{j=0}^{n}.$
@@
$\mathbf{T}^{*}\boldsymbol{\mathbf{v}}_{i}~{}=~{}\hat{\lambda}_{i}\boldsymbol{\mathbf{v}}_{i}$
@@
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct Koopman eigen-functions (Definition 3.8) and the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative (Definition 3.10).
If the matrix pair $(\mathbf{X},\mathbf{Y})$ is linearly consistent (Definition 3.11) and there are at-least as many DMD eigen-values $(n)$ as there are Koopman eigen-values $(r)$, then, the Koopman eigen-values are a subset of the DMD eigen-values:
@@
Given the time-series data $\mathbf{Z}\in\mathbb{C}^{m\times(n+1)}$, the order or capacity of the resulting DMD model is defined thus:
@@ 3.3 Mean subtracted Dynamic Mode Decomposition ($\boldsymbol{\mathbf{\mu}}\textrm{DMD}$)
Mean subtraction is a useful pre-processing step for DMD, when the time series, $\mathbf{Z}$, represents the solution to certain partial differential equations (PDEs) [7]. In this amendment, elaborated here as Algorithm 2, one simply removes the temporal mean of $\mathbf{Z}$ from each of its’ columns before performing DMD. The upshot of such a pre-processing is that the resulting DMD eigen-values and modes Eq. 11 can generate forecasts that respect the boundary conditions associated with the underlying PDE.
@@
    Reduction to a Discrete Fourier Transform (DFT): The DMD eigen-values are restricted to be the $(n+1)$th roots of unity, modulo 1. As a result, the mean-subtracted DMD modes, $\{\,\tilde{\boldsymbol{\mathbf{d}}_{i}}\,\}_{i=1}^{n}$, coincide with the Fourier modes of the original time-series, $\mathbf{Z}$.
@@ 3.3.1 Towards untangling the equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ to DFT
    $\boldsymbol{\mathbf{\phi}}~{}:=~{}[\phi_{i}]_{i=1}^{r}\implies\exists~{}\tilde{\mathbf{C}}\in\mathbb{C}^{m\times r}~{}\textrm{such that}~{}\boldsymbol{\mathbf{\psi}}~{}=~{}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}.$
@@
    $\forall~{}i,\quad\tilde{\mathbf{C}}\boldsymbol{\mathbf{e}}_{i}\neq\boldsymbol{\mathbf{0}}.$
@@
    Under-sampled $(n<r)$: The DMD model has insufficient capacity to describe the $r$-term Koopman mode expansion.
@@
      Just-sampled $(n=r)$: The DMD model has the exact capacity needed to capture the $r$ distinct Koopman eigen-values.
@@
      Over-sampled $(n\geq r+1)$: The capacity of the DMD model is more than sufficient to represent all $r$ Koopman eigen-values.
@@ 4 Exploring the what and where of DMD-DFT equivalence
We begin by developing a necessary and sufficient condition for the equivalence of mean-subtracted DMD and DFT (Theorem 4.1). Building on this result, we find that Lemma 3.6 is a stricter sufficient condition for DMD-DFT equivalence and show that its converse does not hold (Theorem 4.3). Then, we study the relationship between mean-subtracted DMD and temporal DFT as a function of the order [^5] See Definition 3.5 aka capacity of the DMD model. We find that the occurrence of DMD-DFT equivalence is, for all practical purposes, dictated by the model capacity, $n$, and the number of Koopman modes, $r$, comprising the DMD observables $\boldsymbol{\mathbf{\psi}}$ (Theorems 4.5 and 4.8). Finally, we demonstrate that the requirements for the above results -linearly consistent data, and to a lesser extent, a Koopman invariant $\boldsymbol{\mathbf{\psi}}$ can be met in practice using time delay embedding (Proposition 4.10).
@@ 4.1 A geometric characterization of DMD-DFT equivalence
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ KEFs with distinct eigen-values $\{\,\lambda_{i}\,\}_{i=1}^{r}$ and the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative. Furthermore, assume that none of these eigen-values takes the value of $1$ and the model order $(n)$ lies between $2$ and $r$:
@@ 4.3.1 Koopman invariance via time delays
Suppose the set of observables $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct Koopman eigen-functions. Consider augmenting $\boldsymbol{\mathbf{\psi}}$ by taking $d$ time delays:
@@ 4.3.2 DMD-DFT equivalence under delay embedding
Porting our findings on DMD-DFT equivalence to this strategy requires that we extend the notions of under, just and over-sampling to $\mathbf{Z}_{\rm d-delayed}$. The hyper-parameter, $n$, which determines the sampling regime in Algorithms 1 and 2, also happens to be the associated model order (Definition 3.5). This suggests that generalizing the sampling regimes to Algorithm 3 should build upon its model order, $\theta$ [^8] Which, by Definitions 3.5 and 15, has a value of $n-d$.. In particular, under-sampling would translate to $\theta<r$, just-sampling would mean $\theta=r$ and over-sampling would denote $\theta>r$. Hence, we can clarify the notion of DMD-DFT equivalence for $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$:
@@ 5 Numerical experiments
The guarantees on DMD-DFT equivalence, in Corollary 4.14, can be computationally illustrated and tested [^9] All code used in this analysis can found at https://github.com/gowtham-ss-ragavan/msub_mdselect_dmd.git, despite being abstract and fragile. The former can be addressed by translating the mathematical statements into concrete computational tasks, and the latter by testing the translated guarantees on an ensemble of trajectories. The resulting framework for numerically probing Corollary 4.14 is first deployed in scenarios where all the attendant conditions are known to be met. Subsequently, we repeat the same experiments in situations where the requisites of Corollary 4.14 may not be met, and draw upon its’ contrapositive to generate lower bounds on the number of Koopman modes required to represent the observations.
@@ 5.1 Preamble: Translation and trial-by-ensemble
5.1.1 A computation-friendly recast of Corollary 4.14
Validating Corollary 4.14 requires that we assess its’ predictions about the presence or absence of DMD-DFT equivalence, over an appropriate range of model orders. Alas, by Definition 3.7, the associated predictions take the form of set (in)equalities- properties that cannot be reliably discerned when the underlying computations use finite precision arithmetic. So, we leverage Theorem 4.1 to re-phrase those predictions in a numerically tangible fashion, in terms of a real-valued indicator named Relative distance to DFT [^10] For a precise construction, see Definition D.32. This numerical quantity takes values on the interval $[0,1]$, with 00 indicating DMD-DFT equivalence and any other value in $(0,1]$ representing non-equivalence. Consequently, we may recast Corollary 4.14 thus:
@@
5.1.2 Addressing a persistent fragility with ensemble studies
A non-zero value of Relative distance to DFT is useless unless it can be distinguished from 00. Even if we do see this happen for a given trajectory, such an observation does not preclude the possibility of poor discernibility for other trajectories.
Consequently, we systematically test Corollary 5.1 for an ensemble of trajectories. For any choice of a dynamical system Eq. 1, we pick an appropriate range of Companion-orders and #$\#[delays]$ in which Corollary 5.1 is to be examined. For each choice of model-order $(\theta)$ and delay embedding dimension $(d)$, we compute Relative distance to DFT over an ensemble of trajectories and dictionaries.
To see this in detail, suppose $\theta_{\max}$ is the maximal Companion-order and $d_{\max}$ the largest number of time delays to be taken. Consider the trajectory of length $\theta_{\max}+d_{\max}+1$ beginning at $\boldsymbol{\mathbf{\upsilon}}_{1}$ and observed through the vector of $m$ observables $\boldsymbol{\mathbf{\psi}}$. The length of the trajectory has been chosen so that we can compute the two numerical indicators for every pair $(\theta,d)$ in the entire parameter range. For the choice of $(\theta,d)$, the time series matrix $\mathbf{Z}$ reads thus:

### v6

@@ Abstract
Pre-processing time-sequential training data by removing the temporal mean has been a point of contention in the Companion matrix formulation of DMD. This stems from the potential of said pre-processing to render DMD equivalent to a temporal Discrete Fourier Transform (DFT). We prove that this equivalence is impossible when the training data is linearly consistent and the order of the DMD model exceeds the number of Koopman modes. Since model order and training set size are synonymous in this variant of DMD, the parity of DMD and DFT can, therefore, be indicative of inadequate training data.
@@ 3.1 The Koopman operator
$=~{}\alpha_{1}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\psi_{1}\circ\Gamma}~{}+~{}\alpha_{2}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\psi_{2}\circ\Gamma}$
@@ 3.2 Dynamic Mode Decomposition
Dynamic Mode Decomposition [21] is an algorithm that takes time-sequential observations and returns estimates of the constituent Koopman eigen-values and modes. The input to DMD is generated by first choosing $m$ observables of interest (collectively referred to as the dictionary), $\{\psi_{i}\}_{i=1}^{m}$, evaluating them along an orbit, $\{\Gamma^{j-1}\left(\boldsymbol{\mathbf{\upsilon}}_{1}\right)\}_{j=1}^{n+1}$, and collecting the resulting observations (aka snapshots) as the columns of a matrix $\mathbf{Z}$:
(4) $\boldsymbol{\mathbf{\psi}}:=[\psi_{i}]_{i=1}^{m},\quad\mathbf{Z}~{}:=~{}\left[\boldsymbol{\mathbf{\psi}}\left(\Gamma^{j-1}\left(\boldsymbol{\mathbf{\upsilon}}_{1}\right)\right)\right]_{j=1}^{n+1}.$
@@
(8) $\mathbf{T}^{*}\boldsymbol{\mathbf{v}}_{i}~{}=~{}\hat{\lambda}_{i}\boldsymbol{\mathbf{v}}_{i}$
@@
If the matrix pair $(\mathbf{X},\mathbf{Y})$ is linearly consistent and there are at-least as many DMD eigenvalues $(n)$ as there are Koopman eigenvalues $(r)$, then, the Koopman eigenvalues are a subset of the DMD eigen-values:
@@
Given the time-series data $\mathbf{Z}\in\mathbb{C}^{m\times(n+1)}$, the order of the resulting DMD model is defined thus:
@@ 3.3 Mean subtracted Dynamic Mode Decomposition ($\boldsymbol{\mathbf{\mu}}\textrm{DMD}$)
Mean subtraction is a useful pre-processing step for DMD, when the time series, $\mathbf{Z}$, represents the solution to certain partial differential equations (PDEs) [7]. In this amendment, elaborated here as Algorithm 2, one simply removes the temporal mean of $\mathbf{Z}$ from each of its’ columns before performing DMD.
@@
Such a pre-processing enables Koopman eigen-functions with eigenvalue $1$ to be exactly forecast.
Definition 3.6 (Forecasting with $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$).
Let $\mathbf{Z}_{\rm test}$ denote the time series generated by observing the trajectory beginning at $\boldsymbol{\mathbf{\upsilon}}_{\rm test}$ for $n$ time steps:
(13) $\mathbf{Z}_{\rm test}~{}:=~{}\left[\boldsymbol{\mathbf{\psi}}\left(\Gamma^{j-1}\left(\boldsymbol{\mathbf{\upsilon}}_{\rm test}\right)\right)\right]_{j=1}^{n+1}.$
In other words, the $j$-th column of $\mathbf{Z}_{\rm test}$, denoted $(\boldsymbol{\mathbf{z}}_{\rm test})_{j}$, is defined thus:
(14) $(\boldsymbol{\mathbf{z}}_{\rm test})_{j}~{}:=~{}\boldsymbol{\mathbf{\psi}}\left(\Gamma^{j-1}\left(\boldsymbol{\mathbf{\upsilon}}_{\rm test}\right)\right).$
Additionally, let $\boldsymbol{\mathbf{\mu}}_{\rm test}$ denote the temporal mean of $\mathbf{Z}_{\rm test}$ i.e.,
(15) $\boldsymbol{\mathbf{\mu}}_{\rm test}~{}:=~{}\frac{1}{n+1}\,\sum_{j=1}^{n+1}(\boldsymbol{\mathbf{z}}_{\rm test})_{j}.$
Then, $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ forecasts the $j$-th column of $\mathbf{Z}_{\rm test}$ thus:
(16) $(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{j}~{}:=~{}\\ \begin{cases}(\boldsymbol{\mathbf{z}}_{\rm test})_{j},\quad&j~{}=~{}1,\dots,n.\\ \\ \boldsymbol{\mathbf{\mu}}_{\rm test}~{}+~{}\begin{bmatrix}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{j-n+k}-\boldsymbol{\mathbf{\mu}}_{\rm test}\end{bmatrix}_{k=0}^{n-1}~{}~{}\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}],\quad&j~{}>~{}n.\end{cases}$
Theorem 3.7 ($\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ perfectly forecasts constant KEFs).
Suppose we have a vector of observables, $\mathbf{E}^{H}\boldsymbol{\mathbf{\psi}}$, that remains constant along every trajectory i.e.,
(17) $\forall~{}(\boldsymbol{\mathbf{\upsilon}}_{\rm test},j),\quad\mathbf{E}^{H}(\mathbf{z}_{\rm test})_{j}~{}=~{}\mathbf{E}^{H}(\mathbf{z}_{\rm test})_{1}.$
Then, this property is respected by the $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ forecasts as well:
(18) $\forall~{}(\boldsymbol{\mathbf{\upsilon}}_{\rm test},j),\quad\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{j}~{}=~{}\mathbf{E}^{H}(\tilde{\boldsymbol{\mathbf{z}}}_{\rm test})_{1}.$
Proof 3.8.
Refer Appendix A.
Consequently, when the underlying dynamical system Eq. 1 is derived from a PDE, the associated boundary conditions are always satisfied by the $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ forecasts.
@@
    Reduction to a Discrete Fourier Transform (DFT): The DMD eigen-values are restricted to be the $(n+1)$th roots of unity, modulo 1. As a result, the mean-subtracted DMD modes, $\{\,\tilde{\boldsymbol{\mathbf{d}}_{i}}\,\}_{i=1}^{n}$, coincide with the Fourier modes of the original time-series, $\mathbf{Z}$. More specifically, if $\omega$ denotes the $(n+1)$-th root of unity,
    $\omega~{}:=~{}\mathrm{e}^{\mathrm{i}\left(\frac{2\pi}{n+1}\right)},$
    and $\tilde{\boldsymbol{\mathbf{f}}}_{i}$ represents the $i$-th Fourier mode of $\mathbf{Z}$,
    $\tilde{\boldsymbol{\mathbf{f}}}_{i}~{}:=~{}\frac{1}{n+1}\,\sum_{j=1}^{n+1}\boldsymbol{\mathbf{z}}_{j}\,\omega^{-i(j-1)},$
    then, we have
    $\tilde{\lambda}_{i}~{}=~{}\omega^{i}~{}~{}\&~{}~{}\tilde{\boldsymbol{\mathbf{d}}_{i}}~{}=~{}\tilde{\boldsymbol{\mathbf{f}}}_{i},\quad\forall\quad i~{}\in~{}\{1,\ldots,n\}.$
@@ 3.3.1 Towards untangling the equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ to DFT
    (20) $\boldsymbol{\mathbf{\phi}}~{}:=~{}[\phi_{i}]_{i=1}^{r}\implies\exists~{}\tilde{\mathbf{C}}\in\mathbb{C}^{m\times r}~{}\textrm{such that}~{}\boldsymbol{\mathbf{\psi}}~{}=~{}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}.$
@@
    (21) $\forall~{}i,\quad\tilde{\mathbf{C}}\boldsymbol{\mathbf{e}}_{i}\neq\boldsymbol{\mathbf{0}}.$
@@
    Under-sampled $(n<r)$: The order of the DMD model is inadequate for describing the $r$-term Koopman mode expansion.
@@
      Just-sampled $(n=r)$: The DMD model has the exact order needed to capture the $r$ distinct Koopman eigen-values.
@@
      Over-sampled $(n\geq r+1)$: The order of the DMD model is more than sufficient to represent all $r$ Koopman eigen-values.
@@ 4 Exploring the what and where of DMD-DFT equivalence
We begin by developing a necessary and sufficient condition for the equivalence of mean-subtracted DMD and DFT (Theorem 4.1). Building on this result, we find that Lemma 3.9 is a stricter sufficient condition for DMD-DFT equivalence and show that its converse does not hold (Theorem 4.3). Then, we study the relationship between mean-subtracted DMD and temporal DFT as a function of the DMD model order [^5] See Definition 3.5. We find that the occurrence of DMD-DFT equivalence is, for all practical purposes, dictated by the model order, $n$, and the number of Koopman modes, $r$, comprising the DMD observables $\boldsymbol{\mathbf{\psi}}$ (Theorems 4.5 and 4.8). Finally, we demonstrate that the requirements for the above results -linearly consistent data, and to a lesser extent, a Koopman invariant $\boldsymbol{\mathbf{\psi}}$ can be met in practice using time delay embedding (Proposition 4.10).
@@ 4.1 A geometric characterization of DMD-DFT equivalence
Suppose we have $r$ KEFs, $\{\,\phi_{i}\,\}_{i=1}^{r}$, corresponding to distinct eigen-values $\{\,\lambda_{i}\,\}_{i=1}^{r}$. Let the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ be spectrally informative about these $r$ KEFs. Furthermore, assume that none of these eigen-values takes the value of $1$ and the model order $(n)$ lies between $2$ and $r$:
@@ 4.3.1 Koopman invariance via time delays
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs. Consider augmenting $\boldsymbol{\mathbf{\psi}}$ by taking $d$ time delays:
@@ 4.3.2 DMD-DFT equivalence under delay embedding
Porting our findings on DMD-DFT equivalence to this strategy requires that we extend the notions of under, just and over-sampling to $\mathbf{Z}_{\rm d-delayed}$. The hyper-parameter, $n$, which determines the sampling regime in Algorithms 1 and 2, also happens to be the associated model order (Definition 3.5). This suggests that generalizing the sampling regimes to Algorithm 3 should build upon its model order, $\theta$ [^8] Which, by Definitions 3.5 and 24, has a value of $n-d$.. In particular, under-sampling would translate to $\theta<r$, just-sampling would mean $\theta=r$ and over-sampling would denote $\theta>r$. Hence, we can clarify the notion of DMD-DFT equivalence for $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$:
@@ 5 Numerical experiments
The guarantees on DMD-DFT equivalence, in Corollary 4.14, can be computationally illustrated and tested [^9] All code used in this analysis can found at https://github.com/gowtham-ss-ragavan/msub_mdselect_dmd.git, despite being abstract and fragile. The former drawback can be addressed by translating the mathematical statements into concrete computational tasks, and the latter by testing the translated guarantees on an ensemble of trajectories (Section 5.1). The resulting framework for numerically probing Corollary 4.14 is first deployed in scenarios where all the attendant conditions are known to be met (Section 5.2). Subsequently, we repeat the same experiments in situations where the requisites of Corollary 4.14 may not be met, and draw upon its’ contrapositive to generate lower bounds on the number of Koopman modes required to represent the observations (Section 5.3).
@@ 5.1 Preamble: Translation and trial-by-ensemble
5.1.1 A computation-friendly reformulation of Corollary 4.14
Validating Corollary 4.14 requires that we assess its’ predictions about the presence or absence of DMD-DFT equivalence, over an appropriate range of model orders. Alas, by Definition 4.13, the associated predictions take the form of set (in)equalities- properties that cannot be reliably discerned when the underlying computations use finite precision arithmetic.
So, we re-phrase those predictions using Theorem 4.1 in terms of a real-valued and numerically tangible indicator named Relative distance to DFT. We begin by rephrasing Theorem 4.1 in the context of $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$.
Lemma 5.1 (Recasting $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}\equiv\textrm{DFT}$ as a vector equality).
Let $\boldsymbol{\mathbf{\mu}}_{\rm d-delayed}$ denote the temporal mean of $\mathbf{Z}_{\rm d-delayed}$,
$\boldsymbol{\mathbf{\mu}}_{\rm d-delayed}:=\mathbf{Z}_{\rm d-delayed}\frac{\boldsymbol{\mathbf{1}}_{\theta+1}}{\theta+1},$
and $(\mathbf{Z}_{\rm d-delayed})_{\rm ms}$ represent the mean-subtracted version of $\mathbf{Z}_{\rm d-delayed}$:
Additionally, let $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$ be the sub-matrix corresponding to the first $\theta$ columns of $(\mathbf{Z}_{\rm d-delayed})_{\rm ms}$:
Then, a necessary and sufficient condition for delay-embedded $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ to be equivalent to a temporal DFT is that the constant vector, $\boldsymbol{\mathbf{1}}_{\theta}$, be orthogonal to the null-space of $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$:
$d\boldsymbol{\mathbf{\mu}}\textrm{DMD}\equiv\textrm{DFT}~{}~{}\iff~{}~{}\mathcal{P}_{\mathcal{N}\left(\,(\mathbf{X}_{\rm d-delayed})_{\rm ms}\,\right)}\left(\boldsymbol{\mathbf{1}}_{\theta}\right)~{}=~{}\boldsymbol{\mathbf{0}}.$
As such, the projection of $\boldsymbol{\mathbf{1}}_{\theta}$ onto the null-space of $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$ can measure the deviation of $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ from DFT:
Definition 5.2 (Relative distance to DFT).
(28) $\texttt{Relative distance to DFT}\,[\theta,d]:=\frac{\lVert\mathcal{P}_{\mathcal{N}\left(\,(\mathbf{X}_{\rm d-delayed})_{\rm ms}\,\right)}\left(\boldsymbol{\mathbf{1}}_{\theta}\right)\rVert}{\sqrt{\theta}}.$
Therefore, Relative distance to DFT takes values on the interval $[0,1]$, with 00 indicating DMD-DFT equivalence and any other value in $(0,1]$ representing non-equivalence.
Consequently, we may recast Corollary 4.14 thus:
@@
Remark 5.4.
In computing Relative distance to DFT, we do not use the matrix $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$ directly as it becomes ill-conditioned for large values of $d$.
Instead, we work with $(\hat{\mathbf{X}}_{\rm d-delayed})_{\rm ms}$ which is the low rank approximation of $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$ obtained by retaining as many leading singular vectors as possible without the condition number of the approximant exceeding $10^{8}$. Specifically, if we denote the best $s$-rank approximation of $(\mathbf{X}_{\rm d-delayed})_{\rm ms}$ as $\left\llbracket(\mathbf{X}_{\rm d-delayed})_{\rm ms}\right\rrbracket_{s}$, then, $(\hat{\mathbf{X}}_{\rm d-delayed})_{\rm ms}$ is defined as follows:
(29) $s^{*}{}{}{}{}{}{}{}{}{}{}$ $:=~{}\max\left\{s~{}|~{}\textrm{Condition number}\left(\left\llbracket(\mathbf{X}_{\rm d-delayed})_{\rm ms}\right\rrbracket_{s}\right)\leq 10^{8}\right\}$
$(\hat{\mathbf{X}}_{\rm d-delayed})_{\rm ms}{}$ $:=~{}\left\llbracket(\mathbf{X}_{\rm d-delayed})_{\rm ms}\right\rrbracket_{s^{*}}.$
Here, we cap the condition number of the approximation at $10^{8}$ to ensure that when the inputs of $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ are defined up to IEEE double precision of $16$ digits, the $d\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ model, $\boldsymbol{\mathbf{c}}^{*}\left[(\mathbf{Z}_{\rm d-delayed})_{\rm ms}\right]$, has at least $8$ digits of precision [25, 8].
Hence, whenever $\left\lVert(\mathbf{X}_{\rm d-delayed})_{\rm ms}-(\hat{\mathbf{X}}_{\rm d-delayed})_{\rm ms}\right\rVert$ is non-trivial, we end up computing only an upper bound on Relative distance to DFT.
5.1.2 Quantifying a theoretical loophole with ensemble studies
A non-zero value of Relative distance to DFT is useful only if it can be numerically distinguished from 00.
Even if we do see this happen for a given trajectory, such an observation does not preclude the possibility of poor discernibility for other trajectories.
Consequently, we systematically test Corollary 5.3 for an ensemble of trajectories. For any choice of a dynamical system Eq. 1, we pick an appropriate range of model-orders and delay embedding dimensions where Corollary 5.3 is to be examined. For each choice of model-order $(\theta)$ and delay embedding dimension $(d)$, we compute Relative distance to DFT over an ensemble of trajectories and dictionaries.
To see this in detail, suppose $\theta_{\max}$ is the maximal model-order and $d_{\max}$ the largest number of time delays to be taken. Consider the trajectory of length $\theta_{\max}+d_{\max}+1$ beginning at $\boldsymbol{\mathbf{\upsilon}}_{1}$ and observed through the vector of $m$ observables $\boldsymbol{\mathbf{\psi}}$. The length of the trajectory has been chosen so that we can compute Relative distance to DFT  for every pair $(\theta,d)$ in the targeted parameter range. For the choice of $(\theta,d)$, the time series matrix $\mathbf{Z}$ reads thus: